A rectangle has a diagonal of 9, and a width of 5. How would I find the perimeter? Thanks SO much!
Answer by nerdybill(7384)

You have to apply Pythagorean theorem.
The "width, length, and diagonal" of the rectangle forms a "right triangle" enabling you to use the theorem.
.
Let x = length of the rectangle
then
x^2 + 5^2 = 9^2
x^2 + 25 = 81
x^2 = 56
x = 7.483
.
perimeter = 2 (length+ width)
perimeter = 2 (7.483+ 9)
perimeter = 2 (16.483)
perimeter = 32.967